OpenStudy (anonymous):

A rectangular prism has a length of 11 centimeters, a width of 9 centimeters, and a height of 3 centimeters. Its volume is _______ cubic centimeters. What is the cubic centimeters??

jimthompson5910 (jim_thompson5910):

use the formula V = L*W*H V = volume L = length W = width H = height

OpenStudy (anonymous):

so it would be 297. right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

11*9*3 99*3=270+27=297

OpenStudy (anonymous):

ok, thank yous so much!!

jimthompson5910 (jim_thompson5910):

you're welcome